OpenStudy (anonymous):

Can someone help me with this problem? Evaluate the expression, using exact values (use unit circle) sec^2(60deg) + 3cos(210deg)

OpenStudy (tkhunny):

\(\cos(60º) = 1/2\) \(\cos(30º) = \sqrt{3}/2\) These are standard values you should memorize. This is just about all you need to solve the problem.

OpenStudy (anonymous):

I know the values for them I just have a problem putting it in the equation also when you solve it do you ignore the exponent?

OpenStudy (tkhunny):

What's preventing you from doing it? \(\sec(x) = 1/\cos(x)\) Use that and see where it leads.

OpenStudy (anonymous):

The exponent the answer I got for the problem was -3sqrt3

OpenStudy (anonymous):

The exponent is what is confusing me

OpenStudy (tkhunny):

Okay, how did you get that? \(\cos(60º) = 1/2\) \(\sec(60º) = 2\) \(\sec^{2}(60º) = (2)^{2} = 4\) Perhaps your book or teacher failed to tell you that \(\sec^{2}(x) = [\sec(x)]^{2}\)?

OpenStudy (anonymous):

I guess so because I knew I had to use the exponent our teacher just really didn't explain it well to us

OpenStudy (tkhunny):

The notation is a little weird, but it is EVERYWHERE, so we all just have to get used to it. That sort of notation is kind of rare, but you will see it over and over with trigonometric functions.

OpenStudy (anonymous):

How I go the answer was like this.. 2 + 3(-sqrt3/2) 2 - 3sqrt3/2 Then I crossed out the 2's and got -3sqrt3

OpenStudy (tkhunny):

I already showed you how the secant piece is 4. \(3\cos(210º) = -3\cos(30º) = -3(\sqrt{3}/2)\) Thus, \(4 - 3\sqrt{3}/2 = \dfrac{8 - 3\sqrt{3}}{2}\)

OpenStudy (tkhunny):

There is no cancelling of 2s. There is no such thing as cancelling.

OpenStudy (anonymous):

Oh, ok I redid the problem a couple of times and came up with the 8 - 3sqrt3 / 2 answer but I just have one question would you divide the 8 with the 2 to make it 4 - 3sqrt3 as the answer?

OpenStudy (anonymous):

I meant reduce it

OpenStudy (tkhunny):

You must review your Order of Operations. 8 - 3sqrt(3)/2 is NOT correct. [8 - 3sqrt(3)]/2 IS correct. Do you see the difference?

OpenStudy (tkhunny):

4 - 3sqrt(3)/2 = [8 - 3sqrt(3)]/2

OpenStudy (anonymous):

I see the difference you use the brackets to have all the numbers inside be divided by 2

OpenStudy (tkhunny):

Right. Without the brackets, only the numbers/variables next to the 2 are divided by the 2. 4 - 3sqrt(3)/2 = \(4 - \dfrac{3\sqrt{3}}{2}\) [8 - 3sqrt(3)]/2 = \(\dfrac{8 - 3\sqrt{3}}{2}\)

OpenStudy (anonymous):

Ok well thank you for the help :)